# The Chemistry and Physics of 199Hg Nuclear Spin Polarization Relaxation in Quantum Magnetometry Cells

### v1

@@ I Introduction
This research article is different from most because it describes work using optically pumped 199Hg that has been done over the last 40 years, some of which was never published, and with only partial theoretical analysis. As such, a review article is not possible due to the limited specific literature to review. Typically, studies of cell atomic polarization relaxation rates (or inverse of the spin lifetime) were done in the context of building an apparatus to measure some fundamental process of greater immediate interest; as typical in such work, the tedious technical details are barely left as a footnote if presented at all.
@@
The optical pumping of the ground states of 199Hg and 201Hg was first achieved in the early 1960’s by B. Cagnac at the Laboratoire de Spectroscopie Hertzienne de l’ENS (now named the Kastler-Brossel Laboratory). [1] The study of the optical pumping of ground state atoms had been expanding through the 1950’s and Hg was among the very atoms first to achieve spin relaxation times of 100’s of seconds. (See [2, 3] for a review of the field.) Hg is a 1S0 atom, so the ground-state magnetic moment is due only to the nucleus; for 199 $I=1/2$, and for 201, $I=3/2$. Other than application to nuclear magnetic resonance-based gyroscopes [4, 5], optically pumped Hg has found minimal practical applications because alkali atoms, with an unpaired electron spin, provide much greater sensitivities in magnetometer applications.
In the early 1980’s at the University of Washington, Seattle, Prof. E.N. Fortson identified optically pumped Hg as useful for a number of fundamental experiments, the first to search for spatial anisotropy that used the mass quadupole moment of the 201Hg nuclear spin to limit the coupling to its orientation or motion through absolute space in a modernized Hughes-Drever experiment.[6, 7]. The optical pumping cells developed for this work were also used to test a proposed non-linear extension to quantum mechanics, which incidentally provided a gyroscopic measurement of the Earth’s rotation.[8]
@@
However, the main goal of that work was to develop an optical pumping cell in which an electric field could be applied to the processing of 199Hg atoms. If the atom has an electric dipole moment in addition to the magnetic one, there would be a shift in the Larmor frequency and would imply the existence of a fundamental time-reversal violating interaction(s) with the 199 Hg atom, the effects of which are expected to be enhanced in high atomic number atoms. (See [9] for a review of the field.)
@@
The goal at the time (1983-5) was to find a cell configuration that would give a long 199Hg spin polarization lifetime and allow the application of a strong electric field to the atoms, without the need to heat the cell to high temperature and also with a buffer gas (both for high voltage stability). An earlier experiment in the Fortson group employed spin-exchange polarized 129Xe,[10] for an EDM experiment, and lifetimes near room temperature on the order of a thousand or more seconds were routinely achieved in the presence of a field on the order of a few kV/cm. As 199Hg offers a much higher sensitivity to possible fundamental interactions of interest in addition to intrinsically better measurement sensitivity, a vigorous research effort was launched. Attempts to find another spin-1/2 atom to serve as a comagnetometer in the same cell as 199Hg were unsuccessful; however, a workable cell was developed with 199Hg alone, although the spin polarization lifetime would degrade with exposure to 254 nm Hg resonance radiation. This problem persists to this day in that the construction of stable cells appears as hit-and-miss.
The purpose of this paper is to briefly review the development of 199Hg optical pumping cells for EDM and other fundamental experiments and to provide an analysis of data that has not been published. This data sheds light on the chemical and physical mechanisms of lifetime stability, which, when elucidated, will help identify what is needed to attain reproducibility in cell fabrication.
@@
Optically pumped 199Hg has found use in several other fundamental experiments, for example, a tests for axions and their long-range interaction from both spin-polarized and unpolarized material bodies.[11, 12] All of these experiments, including the 199Hg EDM experiments, employ closed cells in that they are filled with appropriate amounts of 199Hg and other gases and permanently sealed off from a preparation manifold.
@@
In addition, 199Hg has served as a comagnetometer for several neutron EDM experiments, the first application being at the Institut Laue-Langevin.[13, 14] In these experiments, the storage cell for ultracold neutrons (UCN) and 199Hg together needed to be developed, and lifetime stability issues persisted in these experiments. A coating that is suitable for both the atomic spin and storage of UCN was found [15], and to date, there does not yet appear to be a fully satisfactory replacement. These experiments are “open" in that background gases (including 199Hg) can be pumped out of the cell.
The spin relaxation lifetime issues for the closed and open cells are related, as will be examined here. There has been continued interest in neutron EDM and specifically experiments that employ a 199Hg comagnetometer.[16] Although the 199Hg EDM limit[17] is numerically much lower than the neutron EDM (this allows it to be used as a comagnetometer), interpretation of the former results in a sensitivity loss regarding fundamental interactions, so the impacts of both types of experiments on testing the standard model of electroweak interactions are similar. New neutron EDM experiments employing 199Hg magnetometry are currently being developed.[18, 19]
@@ I.1 Development of 199Hg EDM cells
Research toward developing a room temperature 199Hg optical pumping cell was begun in 1984, with the first trials using blown fused-silica bulbs of about 18 mm ID. The kinetic mean free path (mfp) between wall collisions is given by
@@
where $V$ and $A$ are the internal volume and surface area, and $R$ is the sphere radius. These cells also had fill stems about 20 mm long where they were sealed from the vacuum system after being filled with natural Hg 0 C vapor pressure.
@@
The first cell had the inner surface coated with paraffin (dotriacontane) and had a spin polarization lifetime of about 40 s. For a larger cell as envisioned for EDM work, with a factor of 2 greater mfp, assuming homogeneous surface relaxation, the lifetime would be 80 s, and good enough for an EDM experiment.
@@
Under illumination with the 254 nm pumping light, the cell lifetime deteriorated. It could be temporarily restored by remelting the wax; overall, ominous effects that we did not fully appreciate at the time.
@@
The next bath of spherical cells had siliconizing agents applied to the inner surface, both Surfasil® and Aquasil® were tested, both with and without wax, and one Aquasil® cell was filled with 40 torr N2 buffer gas (Linde® spectroscopic grade, of unknown history). The wax was introduced only after the cells without wax were removed from the vacuum system. It was observed that the wax did not wet the siliconized surfaces but accumulated in individual droplets, so the wax was driven into the cell stem, where it formed a sort of crystalline plug.
@@
All of these cells had a lifetime of only about 15 sec. This was thought to be due to the damage to the coating that occurred with the cells being sealed off, the poor character of the wax plug, or magnetic field gradients in the case of the buffer gas cell (the magnetic shield was very crude), which had a 5 s. lifetime. The Aquasil® coated cell (10% dilution in acetone, followed by two rinses with acetone and drying by blowing air into the bulb) had the best long-term performance. Warming the cell to 120 $\circ$ C increased useful life to 60 s, with a modest permanent improvement at room temperature.
@@
Emboldened by these results, we made cells similar to those used in [10] but with Hg and only N2 as a buffer gas that is required for high-voltage stability. These cells had a 25 mm OD, 23 mm ID, 10 cm tall fused silica cylinder with a fill stem that enters radially into the side of the cylinder, and capped by fused silica disks of approximately 38mm diameter, glued on with Varian Torr-Seal® epoxy. For Hg, a variety of electrode materials were tried; however, exactly the same electrode coating material (tin oxide) coated with wax with Aquasil® on the cylinder body gave excellent performance.
In Fig. 1, a free-precession signal is shown, obtained by optical pumping with resonance light from a 204Hg microwave discharge lame; there is no hyperfine structure for this isotope, and the single spectral line overlaps the $F=1/2$ to $F=1/2$ level for 199Hg, and the $F=3/2$ to $F=5/2$ for 201Hg (for which no optical pumping signal was ever observed in our EDM cells, as expected).
@@
This plot was obtained by optically pumping the system with circularly polarized resonance light propagating along the direction of a weak magnetic field (10 mG), which was suddenly switched off with the simultaneous sudden application of a second field (10 mG) perpendicular to the light. The angular momentum states for the transition in 199Hg are optimal for optical pumping because the sample becomes transparent (for pure 199Hg) when fully polarized, and the modulation on the light during free precession is 100%, however that modulation is reduced to about 15% with a natural isotopic mixture of Hg.(See Fig. 1 of [2] for a pictorial of this specific case.)
Remarkably, the relaxation time in the dark was about 3600 s. We therefore thought that we were done with cell development, as cells made in this manner proved to be stable under the application of high voltage.
@@
Figure 1: 1st successful EDM cell, probe light limited lifetime to 200 sec, dark lifetime was 3600 s at the time of the measurement.
@@
Two such cells were operated together with electric fields directed oppositely, with the sum of the frequency difference used to stabilize the applied field and a correction signal to stabilize the gradient field and which provided a signal to detect an EDM because the field in the cells was directed oppositely.[20]
The in-phase transmitted light signal provides a measure of the sample atomic polarization, times the atomic density. Shown in Fig. 2 is one of the data runs where the loss of Hg one of the cells is evident as the average transmitted light intensity, measured by a solar blind photomultiplier, steadily increases, representing a loss of Hg atoms (the discharge lamps used in this work were remarkably stable, with few percent at most change in 24 hours). At the same time, the in-phase signal substantially increases, and this is only possible if the wall lifetime increases, which increases the atomic polarization for the approximately constant optical pumping rate.
@@
Figure 2: $V_{DC}$ is a measure of the transmitted light intensity (multiplied by 10 to offset the plot from $V_{in}$), and its increase indicates a loss of Hg atoms. The atomic oscillator in phase signal continues to increase, which is only possible if the cell relaxation rate decreases (e.g. lifetime increases), and then disappears when the Hg density falls toward zero. Data manually digitized from plots in [7], Fig. 8.2, cell 2.
For the case of low absorption, the transmitted light is, by expanding $e^{-x}\approx 1-x$,
$I(t)=I_{0}\left[1-\rho(t)\sigma\ell\big{(}1-\epsilon P(t)\cos\omega t\big{)}\right]$ (2)
where $I_{0}$ is the resonant light intensity, $\rho(t)$ is the time-dependent Hg density, $\sigma$ is the light absorption cross section, $\ell$ is the effective path length for light through the cell, $\epsilon$ is the fraction of light resonant with 199Hg , $P(t)$ is the average magnitude of the atomic polarization, and $\omega$ is the Larmor frequency. Therefore the average $DC$ light intensity, represented by $V_{DC}(t)$, is
$V_{DC}(t)\propto I_{0}\big{(}1-\rho(t)\sigma\ell\big{)}$ (3)
and the in-phase signal is
$V_{in}\propto I_{0}\rho(t)\sigma\ell P(t).$ (4)
The fractional change in the amount of light absorbed by the atoms, relative to the initial, is
$\eta(t)=\frac{V_{DC}(\infty)-V_{DC}(t)}{V_{DC}(\infty)-V_{DC}(0)}$ (5)
$V_{dc}(\infty)\approx V_{dc}(t_{f})+30$ (6)
in the voltage units of Fig. 2, and where $t_{f}=18000$ s is the end of the measurement time however the absorption has not yet reached zero. A rough extrapolation gives a small increase in the absorption.
It is the absorbed light that produces $V_{in}$, so
$\frac{V_{in}(t)}{\eta(t)}\propto P(t).$ (7)
From the basic principles of the atomic oscillator,
$P(t)=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}+\omega_{1}}$ (8)
where $\Gamma_{p}$ is the optical pumping rate, $\Gamma_{w}=1/\tau_{w}$ is the spin relaxation rate and $\omega_{1}$ is the effective precession rate around the oscillating field.
Because the in-phase atomic oscillator signal $V_{in}(t)$ is proportional to the absorbed light $\eta(t)$, it follows that
$\frac{\eta(t)}{V_{in}(t)}\propto(P(t))^{-1}=\frac{\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}}{\Gamma_{p}(t)}.$ (9)
The pumping rate is given by the average DC light intensity in the cell, which is
$\Gamma_{p}\propto[1-\alpha\eta(t)]$ (10)
where $\alpha=0.58$ is the initial light absorption as determined by $V_{dc}(0)$ and $V_{dc}(\infty)$, noting that typically 20% of the light is nonresonant for the microwave discharge lamps we employed.
Therefore,
$\frac{[1-\alpha\eta(t)]\eta(t)}{V_{in}(t)}\propto\Gamma_{p}(t)+\Gamma_{w}(t)+\omega_{1}=\Gamma_{tot}(t).$ (11)
We can take the ratio $\Gamma_{tot}(t)/\Gamma_{tot}(0)$ to determine how the total relaxation lifetime evolves, which is shown in Fig. 3 to decrease. This can be compared to the pumping rate $\Gamma_{p}(t)/\Gamma_{p}(0)$, which increases over the same time period, because the light intensity in the cell increases as Hg is lost. The other term in $\Gamma_{tot}(t)$ is $\omega_{1}$ which is constant.
In typical operation of the atomic oscillators, the light intensity and $\omega_{1}$ were kept constant between data runs. Normally we tried to achieve
$\omega_{1}=\Gamma_{p}\approx\Gamma_{w}/2.$ (12)
This relationship can only be satisfied for the data in Fig. 3 if $\Gamma_{w}(0)>5\Gamma_{p}$, a factor 2.5 times larger than the target value. However, this choice provided a reasonable balance over the run, plus stronger pumping tended to degrade the cells faster.
This shows that, without a doubt, during this data run, $\Gamma_{tot}(t)$ decreased with time, so $\tau_{w}$ must have increased as Hg was consumed. From this graph, we estimate that $\tau_{w}$ increased by at least a factor of three over the duration of the experimental run (data near the end of the run are unreliable due to the failure of the atomic oscillator as the signal diminished) to within the approximations and the lack of experimental specifics for work done nearly 40 years ago. However, this is the first clear evidence that 254 nm resonance radiation irradiation can be beneficial in one way, and perhaps the only direct evidence to date.
In principle, a source of Hg in the cell could be along with a small amount of oxygen; however, the optical density would be too high for any practical system at the time.
Figure 3: By use of the cell light transmission, it is possible to deduce the the pump rate $\Gamma_{p}$, and the in-phase signal to deduce $\Gamma_{p}+\Gamma_{w}+\omega_{1}$. As the last term is constant, and their sum is decreasing while $\Gamma_{p}$ is increasing, is clear that the wall relaxation rate is decreasing. The loss of Hg atoms eventually lead to a failure of the atomic oscillator for this cell, near the final time of the plotted data.
In obtaining the first 199Hg EDM limit, new cells were manufactured for each short data run and the result of that was published [7, 20]. .
@@
3P1 Hg 4.89 allowed transition ($f\approx 1$)

### v2

# Surface and chemical effects on ${{}^{1}99}$Hg spin polarization relaxation in optically pumped magnetometers

@@ I Introduction
This research article is different from most because it describes work using optically pumped 199Hg that has been done over the last 40 years, some of which was never published, and with only partial theoretical analysis. As such, a review article is not possible due to the limited specific literature to review. Typically, studies of cell wall-induced atomic polarization relaxation times, defined here as $\tau_{w}$ (with wall relaxation rate $\Gamma_{w}=\tau^{-1}_{w}$), were performed in the context of building an apparatus to measure some fundamental process of greater immediate interest; as typical in such work, tedious technical details are barely left as a footnote if presented at all.
@@
The study of the optical pumping of ground state atoms had been expanding through the 1950s and Hg was among the very atoms first to achieve $\tau_{w}$ of 100’s of seconds. One of the most important developments in the field is transmission monitoring of the ensemble spin polarization, an idea due to Dehmelt [1] who patented a magnetometer design that is still in use.[2] (See [3, 4] for reviews of the field.)
The optical pumping of the ground states of 199Hg and 201Hg was first achieved in the early 1960s by B. Cagnac at the Laboratoire de Spectroscopie Hertzienne de l’ENS (now named the Kastler-Brossel Laboratory).[5] The theory was fully developed and tested by Cohen-Tannoudji in his Ph.D. studies[6, 7], and he has written a general review of the early work at this laboratory. [8]
Hg is a 1S0 atom, so the ground-state magnetic moment, if any, is due only to the nucleus; for the two isotopes with significant natural abundance, only two have non-zero nuclear spin: 199 $I=1/2$, and for 201, $I=3/2$. In addition to applications to nuclear magnetic resonance-based gyroscopes [9, 10], optically pumped Hg has found minimal practical applications because alkali atoms, with an unpaired electron spin and hence larger atomic gyromagnetic ratios, can provide much higher sensitivities in magnetometer applications. Optically pumped Hg has, however, found application in several fundamental experiments.
In the early 1980s at the University of Washington, Seattle, Prof. E.N. Fortson identified optically pumped Hg as useful for a number of fundamental experiments, the first to search for spatial anisotropy that used the mass quadupole moment of the 201Hg nuclear spin to limit the coupling to its orientation or motion through absolute space in a modernized Hughes-Drever experiment.[11, 12]. The optical pumping cells developed for this work were also used to test a proposed non-linear extension to quantum mechanics, which incidentally provided a gyroscopic measurement of the Earth’s rotation.[13]
@@
These experiments were side projects; the main research goal at the time (1983-90) was to develop an optical pumping cell in which an electric field could be generated inside the cell volume and to determine whether the applied electric field leads to a change in the Larmor precession frequency of the nuclear spin of 199Hg atoms. If the atom has an electric dipole moment in addition to the magnetic one, there will be a shift in the Larmor frequency and this will imply the existence of a fundamental time-reversal violating interaction(s) with the 199 Hg atom, the effects of which are expected to be enhanced in high atomic number atoms. (See [14] for a review of the field, and a more recent review by Chupp et al. [15].)
@@
The essential problem was to find cell coating materials that give a long $\tau_{w}$ for 199Hg and also allow the application of high voltage, without the need to heat the cell to high temperature. Inclusion of a relatively inert “buffer gas together with low (room) temperatures is required for high voltage stability. An earlier experiment in the Fortson group employed spin-exchange polarized 129Xe,[16] for an EDM experiment, and $\tau_{w}$ near room temperature on the order of a thousand or more seconds were routinely achieved in the presence of a field on the order of a few kV/cm. As 199Hg offers a higher sensitivity to possible fundamental interactions of interest in addition to an intrinsically better measurement sensitivity, a vigorous research effort was launched to establish an 199Hg EDM experiment.
Attempts to find another spin-1/2 atom to serve as a comagnetometer in the same cell as 199Hg were unsuccessful; however, a workable cell was developed with 199Hg alone, although $\tau_{w}$ would degrade with exposure to 254 nm Hg resonance radiation. (201Hg has a very short $\tau_{w}$ in room-temperature cells due to its nuclear quadrupole moment, which in addition, anyway, precludes its use in EDM experiments.) This problem persists to this day in that the construction of stable cells remains a hit-and-miss.
@@
Optically pumped 199Hg has found use in several other fundamental experiments, for example, a tests for axions and their long-range interaction from both spin-polarized and unpolarized material bodies.[17, 18] All of these experiments, including the 199Hg EDM experiments, employ closed cells in that they are filled with appropriate amounts of 199Hg vapor and other gases and permanently sealed off from a preparation manifold.
@@
In addition, 199Hg has served as a comagnetometer for several neutron EDM experiments, the first application being at the Institut Laue-Langevin.[19, 20] In these experiments, the storage cell for ultra-cold neutrons (UCN) and 199 Hg together needed to be developed, and $\tau_{w}$ stability problems persisted in these experiments. A coating that is suitable for both the atomic spin and storage of UCN was found, deuterated polystrerne (DPS) [21], and to date there does not appear to be a fully satisfactory replacement. These experiments are “open" in that background gases (including 199Hg) can be pumped out of the cell. The techniques were greatly refined by the Paul-Scherrer Institute, [22, 23] leading to the most sensitive neutron EDM measurement performed so far. This apparatus was also used to search for axion coupling between a Cu mass and nuclear spins.[24]
Better wall coatings for 199Hg with longer and stable ensemble nuclear spin polarization relaxation times $\tau_{w}$ are of interest to several ongoing experiments, including the neutron EDM experiments at the Paul-Scherrer Institut [25] and Los Alamos [26], and tests of the equivalence principle by Hunter et al. [27] which is the principal motivation for the study presented here.
The $\tau_{w}$ issues for closed and open cells are related, as will be examined here. The purpose of this paper is to briefly review the development of 199Hg optical pumping cells for EDM and other fundamental experiments and to provide an analysis of data that has not been published. This data sheds light on the chemical and physical mechanisms of the $\tau_{w}$ instability, which, when elucidated, will help identify what is needed to achieve reproducibility in cell fabrication.
As an aside, although the limit of the 199Hg EDM [28] is numerically much lower than the neutron EDM (this allows it to be used as a comagnetometer), theoretical interpretation of the limits of the 199Hg EDM results in a loss of sensitivity to fundamental interactions, so the impacts of both types of experiments on testing the standard model of electroweak interactions are similar. EDM experiments based on optically pumped 199Hg alone have potential for improvement, although the lack of a comagnetometer within the cell volume remains a potential ultimate limitation.
@@ I.1 Development of 199Hg EDM cells
The basic optical pumping technique used in the development of 199Hg cells employed circularly polarized resonance radiation and transmission monitoring [1] of the ensemble spin polarization. Resonance light from a 204Hg microwave discharge lamp contains a single 254 nm line (corresponding to the $6^{1}S_{0}$ to $6^{3}P_{1}$ transition) because there is no hyperfine structure for this isotope.[5, 7] The single spectral line overlaps the $F=1/2$ ground state to $F=1/2$ excited state levels of 199Hg. If a cell containing atoms is placed in a few mG magnetic field oriented along the light propagation direction, the ground state atoms will be excited only from the $m_{F}=-1/2$ if the light has left-handed circular polarization (which carries $+1$ unit of angular momentum). The decay back to the ground state can go to either level; however, atoms accumulate in the $m_{F}=1/2$ level because these never go through an optical pumping cycle. Eventually, the atomic nuclear spin polarization reaches equilibrium, and the transparency of the cell increases. The equilibrium spin polarization is, when the only spin polarization relaxation is due to wall interactions, is
$P=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}}$ (1)
where $\Gamma_{p}$ is the optical pumping rate or the inverse of the pumping cycle time, $1/\tau_{p}$, and $\Gamma_{w}=1/\tau_{w}$ is the wall relaxation rate. For most of the subsequent discussion, it will be assumed that wall relaxation is dominant over magnetic gradient effects or interactions with other atoms or molecules in the gas phase. In the weak field limit ($\ll 100$ G), in the absence of magnetic gradients, the longitudinal and transverse (coherence) relaxation times are equal ($T_{1}$ and $T_{2}$ in nuclear magnetic resonance terminology). (See [3] for a comprehensive review of ground-state optical pumping, in which Fig. 1 is a pictorial of this specific case.)
If the magnetic field is suddenly switched to be perpendicular to the light (in time much faster than the inverse Larmor frequency), $P$ becomes time dependent, and the transmitted light is modulated at the Larmor frequency, which is about 7 Hz for a 10 mG magnetic field. The modulation level will exponentially decay with a total rate given by the sum of $\Gamma_{p}$ and $\Gamma_{w}$, so to determine $\tau_{w}$, the total rate must be extrapolated to $\Gamma_{p}=0$.
For discharge lamps, the non-resonant background light was typically only 20% of the emitted light, with the dominant noise source being the shot noise in the detected photons. Recent experiments have employed frequency-quadrupled semiconductor lasers, which have reduced non-resonant light and higher-efficiency detectors. Furthermore, such lasers are tunable, which allows for a higher density of 199Hg because the laser can be tuned away from the absorption line center.
Research toward developing a room temperature 199Hg optical pumping cell was begun in 1984, with the first trials using fused-silica blown bulbs of about 18 mm ID. The kinetic mean free path (mfp) between wall collisions is given by
@@
where $V$ and $A$ are the internal volume and surface area, and $R$ is the radius of the sphere. These cells also had fill stems about 20 mm long where they were sealed from the vacuum system after being filled with natural Hg 0 C vapor pressure.
@@
The first cell had the inner surface coated with paraffin (dotriacontane) “wax" and had a $\tau_{w}$ of about 40 s. For a larger cell, as envisioned for EDM work, with a factor of 2 greater mfp, assuming homogeneous surface relaxation, $\tau_{w}$ would be 80 s and good enough for an EDM experiment.
@@
Under illumination with the 254 nm pumping light, the cell’s $\tau_{w}$ deteriorated. It could be temporarily restored by remelting the wax; this potentially ominous effect was not fully appreciated at the time.
@@
The next batch of spherical cells had siliconizing agents applied to the inner surface, both Surfasil® and Aquasil® were tested, both with and without wax, and an Aquasil® cell was filled with 40 torr N2 buffer gas (Linde® spectroscopic grade). The wax was introduced only after the cells without wax were removed from the vacuum system. It was observed that the wax did not wet the siliconized surfaces but accumulated in individual droplets, so the wax was driven into the cell stem, where it formed a sort of crystalline plug.
@@
All of these cells had a $\tau_{w}$ of only about 15 s. This was thought to be due to the damage to the coating that occurred with the cells being sealed off, the poor character of the wax plug, or magnetic field gradients in the case of the buffer gas cell (the magnetic shield was very crude), which had a 5 s $\tau_{w}$. The Aquasil® coated cell (10% dilution in acetone, followed by two rinses with acetone and drying by blowing air into the bulb) had the best long-term performance. Warming the cell to 120 C increased useful life to 60 s, with a modest permanent improvement at room temperature.
@@
Emboldened by these results, we made cells similar to those used in the 129Xe experiment [16] but with Hg and only N2 as a buffer gas required for high-voltage stability. These cells had a fused silica cylinder of 25 mm OD, 23 mm ID, 10 mm high with a fill stem that enters radially onto the side of the cylinder, and the cylinder was capped with fused silica disks of approximately 38 mm diameter, glued with Varian Torr-Seal® epoxy. To provide a means of high voltage (HV) application within the cell to establish an electric field, a variety of electrode materials were tried; however, exactly the same electrode coating material (tin oxide) as used in the 129Xe experiment, coated with wax, and with Aquasil® on the cylinder body gave excellent performance.
@@
The data shown in Fig. 1 was obtained by optically pumping the system with circularly polarized resonance light propagating along the direction of a weak magnetic field (10 mG), which was suddenly switched off with the simultaneous sudden application of a second field (10 mG) perpendicular to the light, as described earlier. The angular momentum states for the transition in 199Hg are optimal for optical pumping because the sample becomes transparent (for pure 199Hg) when fully polarized, and the modulation on the light during free precession can be nearly 100% for a fully polarized ensemble; however, the modulation is reduced to about 15% with a natural isotopic mixture of Hg, as was used throughout the cell development period and for the first 199Hg EDM experiment. The actual Larmor frequency was about 10 Hz, which was mixed with the signal from a constant oscillator to bring the precession frequency down to where it could be measured with a chart recorder.
At the minimum controllable light level that we could achieve with our system, $\tau_{w}$ was 300 s, and met the 100 s goal for cell development (because $1/f$ noise magnetic field noise required a relatively fast electric field reversal rate, so a longer time was unnecessary). A quick test in the dark indicated a lifetime in excess of 1000 s, which is consistent with other ${}^{1}S_{0}$ nuclear-spin 1/2 atoms, particularly 129Xe. We therefore thought that we were done with cell development, as cells made in this manner proved to be stable under the application of high voltage.
@@
Figure 1: 1st successful EDM cell, probe light limited the spin polarization relaxation time to 200 s.
@@
Two such cells were operated together with electric fields directed oppositely, with the sum of the frequency difference used to stabilize the average applied field, and the difference used as a correction signal to stabilize the gradient field between the cells. The gradient correction signal was sensitive to an EDM because the electric field in the cells was directed oppositely.[29]
The in-phase oscillating transmitted light signal provides a measure of the sample atomic polarization, times the atomic density. Shown in Fig. 2 is one of the data runs in which the loss of Hg in the cells is evident at the end of the run when the oscillator signal $V_{in}$ fades.
@@
Figure 2: $V_{DC}$ is a measure of the transmitted light intensity (multiplied here by 10 to offset the plot from $V_{in}$), and its increase indicates a general upward drift of the light intensity, with a rapid loss of Hg atoms near the end time, as further evidence by the loss of the oscillator in-phase signal $V_{in}$. This data was manually digitized from plots in [12], Fig. 8.2, cell 2.
This data has some interesting features that can be understood by considering the degree of polarization of the ensemble,
$P=\frac{\Gamma_{p}}{\Gamma_{p}+\Gamma_{w}+\omega_{1}}$ (3)
where $\omega_{1}$ is the oscillating field amplitude. Normally, the light intensity in our experimental work was very stable, so this data run is slightly anomalous; however, given the haste with which we had to set up and test the system, it is understandable. At the start of the experiment, the light intensity and $\omega_{1}$ were adjusted such that $\Gamma_{p}\approx\omega_{1}$ and $\Gamma_{p}+\omega_{1}\approx 50\ {\rm s}^{-1}$, with approximately 30% precision. Noting that the light intensity is proportional to $\Gamma_{p}$, and that $\Gamma_{w}=\tau_{w}^{-1}\ll\Gamma_{p}$, the in-phase signal should scale with the light intensity as its product with $P$:
$V_{in}\propto\frac{\Gamma_{p}^{2}}{\Gamma_{p}+\omega_{1}}\approx\frac{\Gamma_{p}}{2}.$ (4)
This scaling works up to about 13000 seconds, at which time $V_{DC}$ begins to increase rapidly, while $V_{in}$ begins to fall.
This marks the onset of the loss of Hg due to reactions with atomic oxygen, or more accurately, ozone O3 which takes some time to build up. This process will be described in some detail later in this article; is a two-step process by which Hg reacts with O2 to form O3 and atomic O. O3 reacts irreversibly with Hg to form HgO, resulting in an increase in transmitted light. The initial O3 concentration is zero, so the loss of Hg does not begin until it is formed, with the loss rate increasing over irradiation time. The differential equation governing this has a loss rate that increases linearly with time (neglecting the loss of Hg so this is approximate, for the early time development),
$\frac{d\rho}{dt}=-\alpha t\rho$ (5)
where $\rho$ is the Hg density, and $\alpha$ is a constant rate parameter that is proportional to the resonance light intensity. A full model with both varying Hg density and light intensity gives a good representation of the observed time dependence. In the case of constant intensity, the solution is $\rho(t)=\rho(0)e^{-\alpha t^{2}}$ which explains the sudden drop in $V_{in}$ at the end of the run. If a cell had been irradiated, the density would continue to change as $\rho(t)=\rho(t_{0})e^{-\alpha t_{0}t}$, assuming the ozone concentration is constant.
It can be further noted that the change in $V_{DC}$ when the Hg disappears, accounting for the upward drift in light intensity, implies that the Hg light absorption at the beginning of the run was only about 15%. This is most likely due to an earlier loss of Hg from the same photosynthesized reaction. Even if a cell is briefly illuminated, there will be some ozone formation, and although at low concentrations the rate of Hg loss is slow, nonetheless the reactions will continue until the ozone is absorbed or reacted with the materials within the cell, as described above. Note that these cells were operated in the strong pumping regime, so it is not possible to deduce if there was a change in $\tau_{w}$ based on these data.
The loss of Hg limited the duration of the EDM measurements for each cell pair, which for the case discussed was about 5 h. In obtaining the first 199Hg EDM limit, we opted to manufacture new cells (in batches of eight) that would be used in pairs until their Hg density decreased to a point where the oscillators failed. The net result of this imperfect but successful study was published.[12, 29]
In principle, a source of Hg could have been incorporated into cells and thus keep the Hg density constant; however, the optical density would have been too high for any practical system at the time, as the proper natural Hg density roughly corresponds to the $0^{\circ}$ vapor pressure. (199Hg accounts for only 15% of the total absorption of light produced by 204Hg; later cells employed enriched 199Hg .) At the time, no tunable sources of resonance light were available. This allows probing away from the line center, as is done currently using quadrupled semiconductor lasers.
@@
3P1 Hg 4.89 allowed transition (oscillator strength $f\approx 1$)
@@
(H5C6)-H 4.46 extra binding energy compared to C-H due to resonance
stabilization of the phenyl ring